# TRACED: Transition-aware Regret Approximation with Co-learnability for Environment Design

### v2

@@ Abstract
Generalizing deep reinforcement learning agents to unseen environments remains a significant challenge. One promising solution is Unsupervised Environment Design (UED), a co-evolutionary framework in which a teacher adaptively generates tasks with high learning potential, while a student learns a robust policy from this evolving curriculum. Existing UED methods typically measure learning potential via regret, the gap between optimal and current performance, approximated solely by value-function loss. Building on these approaches, we introduce the transition prediction error as an additional term in our regret approximation. To capture how training on one task affects performance on others, we further propose a lightweight metric called co-learnability. By combining these two measures, we present Transition-aware Regret Approximation with Co-learnability for Environment Design (TRACED). Empirical evaluations show that TRACED yields curricula that improve zero-shot generalization across multiple benchmarks while requiring up to 2$\times$ fewer environment interactions than strong baselines. Ablation studies confirm that the transition prediction error drives rapid complexity ramp-up and that co-learnability delivers additional gains when paired with the transition prediction error. These results demonstrate how refined regret approximation and explicit modeling of task relationships can be leveraged for sample-efficient curriculum design in UED.
@@ 1 Introduction
Deep reinforcement learning (RL) has achieved remarkable success in games, continuous control, and robotics [1]. Ideally, we want agents that generalize robustly to a broad range of unseen environments. However, hand-crafting a training distribution that captures all real-world variability is intractable, and agents often overfit even large training sets, performing poorly out-of-distribution [2, 3].
@@
Unsupervised Environment Design (UED) tackles this by adapting the curriculum: a teacher module generates training tasks that challenge the student agent [4]. A popular class of UED methods measures task difficulty by regret, the difference between the optimal return and the agent’s achieved return, and uses this metric to guide curriculum design [4, 5, 6, 7, 8, 9, 10]. Unfortunately, genuine regret requires knowing each environment’s optimal $Q^{*}$, which is infeasible in complex domains. Existing approaches, therefore, resort to coarse proxies such as Positive Value Loss (PVL) or maximum observed return (MaxMC) [11]. In this paper, we refine regret estimation by explicitly decomposing it into (i) a value error, captured by PVL, and (ii) a transition prediction error, estimated via a transition dynamics model.
@@
We further introduce a metric called Co-Learnability to quantify how training on one task benefits others. For instance, consider three 100-word vocabulary tasks, Spanish, English, and Japanese, whose transfer patterns differ: because Spanish and English share many cognates, learning Spanish accelerates lexical access and boosts English accuracy [12, 13, 14], reflecting high Co-Learnability; in contrast, Japanese is typologically distant [15], so gains from Japanese may transfer unefficiently to English, reflecting low Co-Learnability. We present a lightweight estimator of Co-Learnability that leverages observed changes in approximated regret, avoiding any additional modeling overhead in the UED loop.
@@
Figure 1: Task Priority Landscape. Task in the top-right quadrant (high difficulty and high co-learnability) are scheduled with the highest priority in the curriculum.
@@
As illustrated in Figure 1, these two ingredients, refined regret and Co-Learnability, are fused into a unified Task Priority score via a rank-based transform (cf. [7]), which governs both new task generation and prioritized replay. We call our algorithm TRACED (Transition-aware Regret Approximation with Co-learnability for Environment Design). We evaluate TRACED on procedurally generated MiniGrid (MG) [16] and BipedalWalker (BW) [17]. In MG, we compare against Domain Randomization (DR) [18], PLR [5], ADD [19], and ACCEL [7], the strongest of these baselines. In BW, we additionally include the state-of-the-art (sota) CENIE [20].
@@
On 12 held-out MG mazes, TRACED achieves a mean solved rate of 66% after 10k updates, outperforming ACCEL’s 60% after 20k updates. In the extreme maze variants PerfectMaze, TRACED reaches 27% success on PerfectMazeLarge and 10% on PerfectMazeXL, representing 35-40% relative gains over ACCEL with only half the training time. In BW, TRACED achieves an 18% improvement over CENIE while halving the wall-clock time. Ablation studies reveal that the transition-prediction error term drives rapid complexity ramp-up, and that Co-Learnability provides additional gains when combined with our regret term. Collectively, these results pave a new path toward sample-efficient curriculum design in UED by combining refined regret approximation with explicit modeling of task relationships.
@@ 2.1 Underspecified Partially Observable MDPs (UPOMDPs)
We model our environments as underspecified partially observable Markov decision processes (UPOMDPs) following Dennis et al. [4]. A UPOMDP is a tuple $\mathcal{M}=\langle A,O,\Theta,S,P_{0},P_{T},\mathcal{I},\mathcal{R},\gamma\rangle$, where $A$ is a finite set of actions, $S$ is a latent state space, and $O$ is an observation space. The observation function $\mathcal{I}:S\to\Delta(O)$ generates each observation $o_{t}$ given the true state $s_{t}$, the reward function $\mathcal{R}:S\times A\to\mathbb{R}$ and discount factor $\gamma\in[0,1)$ are shared across all levels. Crucially, $\Theta$ is a set of underspecified parameters that distinguish individual “levels”: for each $\theta\in\Theta$, the initial state is drawn from $P_{0}(\theta)\in\Delta(S)$ and transitions follow $P_{T}(s_{t+1}\mid s_{t},a_{t},\theta)$.
At each time step $t$, the agent observes $o_{t}\sim\mathcal{I}(s_{t})$ and selects an action $a_{t}$ according to a trajectory-conditioned policy $\pi\bigl{(}a_{t}|o_{0},a_{0},\dots,o_{t}\bigr{)}$. For a fixed level $\theta$, the utility of policy $\pi$ is the expected discounted return $U_{\theta}(\pi)=\mathbb{E}\!\Bigl{[}\sum_{t=0}^{T}\gamma^{t}\,r_{t}\Bigr{]},r_{t}=\mathcal{R}(s_{t},a_{t})$ with the expectation taken over both the stochastic dynamics and the policy’s choices. We denote an optimal policy on level $\theta$ by $\pi^{\star}_{\theta}\in\arg\max_{\pi}U_{\theta}(\pi)$.
@@ 2.2 Unsupervised Environment Design (UED)
UED provides a series of levels with unknown parameters that are used to produce task environments automatically as a curriculum for the agent so as to efficiently train a single generalist policy $\pi_{\phi}$ across the entire parameter space $\Theta$ [4]. Recent UED methods maximize regret of the agent to generate a distribution of environments that guide effective learning [5, 7, 20]. The regret of policy $\pi$ is defined as the difference in expected reward between optimal policy $\pi^{*}$ and policy $\pi$, $Regret_{\theta}(\pi)=U_{\theta}(\pi^{*})-U_{\theta}(\pi)$.
Because the true optimal policy $\pi^{\star}$ is unknown in complex environments, existing methods approximate the instantaneous regret using proxy metrics. For example, PLR [6] evaluates two such proxies: Positive Value Loss (PVL) and Maximum Monte Carlo (MaxMC). PVL estimates the instantaneous regret as the average positive part of the Generalized Advantage Estimation (GAE)-based Temporal Difference (TD) errors over an episode. PVL for an episode $\tau$ of length $T$ is defined as
@@
where $\gamma$ is the discount factor, $\lambda$ is the GAE coefficient and $\delta_{t}=r_{t}+\gamma\,V(s_{t+1})-V(s_{t})$ is the one-step TD-error at timestep $t$. MaxMC uses the highest undiscounted return observed on the task instead of a bootstrap target. Other criteria such as policy entropy, one-step TD error, GAE, policy min-margin, and policy least-confidence have also been evaluated in curriculum learning contexts [5].
3 Method
At each step, curriculum design aims to select the tasks that will most effectively drive the agent’s learning progress. Our approach enhances UED by combining (1) a refined regret approximation integrating both value and transition errors and (2) a simple yet principled measure of task co-learnability. We compute a unified Task Priority score from these two quantities, which then governs both new task generation and replay sampling in the UED loop.
@@ 3.1 Regret Approximation via Transition Prediction Loss
Since regret quantifies the difficulty an agent experiences on a task, improving regret approximation can lead to more accurate task difficulty estimation and thus more effective curricula.
To address this, we decompose one-step regret at state–action pair $(s,a)$ into two interpretable terms:
$\mathrm{Regret}(s,a)$ $=Q^{*}(s,a)-Q(s,a)$ (2)
$=\gamma\Bigl{(}\mathbb{E}_{s_{1}\sim P(\cdot\mid s,a)}\bigl{[}Q^{*}(s_{1},a^{\prime})\bigr{]}-\mathbb{E}_{s_{2}\sim\hat{P}(\cdot\mid s,a)}\bigl{[}Q(s_{2},a^{\prime\prime})\bigr{]}\Bigr{)}$
$=\gamma\Bigl{(}\underbrace{\mathbb{E}_{s_{1}\sim P}\bigl{[}\,Q^{*}(s_{1},a^{\prime})-Q(s_{1},a^{\prime})\,\bigr{]}}_{\text{(i) Value Error}}+\underbrace{\mathbb{E}_{s_{1}\sim P}\bigl{[}Q(s_{1},a^{\prime})\bigr{]}-\mathbb{E}_{s_{2}\sim\hat{P}}\bigl{[}Q(s_{2},a^{\prime\prime})\bigr{]}}_{\text{(ii) Transition Prediction Error}}\Bigr{)}$
Term (i), the value error, measures the expected discrepancy between the true optimal action-value $Q^{*}$ and the agent’s learned $Q$. Term (ii), the transition prediction error, quantifies the bias introduced by replacing the true transition kernel $P$ with an empirical approximation $\hat{P}$, where $\hat{P}(s^{\prime}\!\mid s,a)$ is the conditional distribution estimated from the agent’s collected experience $\{(s,a,s^{\prime})\}$. Accurately estimating both terms is crucial: term (i) reflects the agent’s value-approximation quality, whereas term (ii) captures how model-mismatch in the dynamics affects long-term returns.
Because $Q^{*}$ and $P$ are unknown, we approximate (i) via the Positive Value Loss (PVL) and (ii) via a learned transition dynamics estimator $f_{\phi}$. Concretely:
Definition 1 (Average Transition Prediction Loss). Train a transition dynamics estimator $f_{\phi}$, implemented as a recurrent model, to minimize a one-step reconstruction loss $L_{\mathrm{trans}}(s_{t},a_{t})$ between the observed next state $s_{t+1}$ and the prediction $\hat{s}_{t+1}=f_{\phi}(s_{t},a_{t})$ (details in Appendix F.3). Over an episode $\tau=(s_{0},a_{0},\dots,s_{T})$, define
@@
where $\alpha>0$ balances value versus transition terms (see Appendix B).
By explicitly incorporating both components, our regret approximation more faithfully reflects task difficulty and yields more effective curricula. After computing the approximated regret $\widehat{\mathrm{Regret}}(\tau)$ for each sampled task instance, we append this value to the Task Difficulty Buffer (TDB). Consequently, the number of entries in the TDB corresponding to task $i$ exactly equals the number of times task $i$ has been sampled (see Figure 2 for an overview).
Figure 2: Task Difficulty Calculation Workflow. The Task Difficulty Buffer (TDB) records each task’s history of approximated regret. The agent interacts with sampled tasks to collect episode trajectories, which are stored in the Rollouts Buffer (RB). For each trajectory, we compute the Positive Value Loss (PVL) and the Average Transition-Prediction Loss (ATPL). Their sum produces the updated task difficulty (approximated regret), which is appended to the TDB to refresh each sampled task’s stored difficulty.
@@ 3.2 Task Priority Construction
A central challenge in UED is to decide which tasks to present to the agent at each step. Intuitively, we want to (1) focus on tasks that remain challenging, while (2) exploiting tasks whose training yields transfer to others. To this end, we introduce two complementary quantities, Task Difficulty and Co-Learnability, and combine them into a single Task Priority score.
Definition 2 (Task Difficulty). Let $s_{i}(t)\;=\;\max\{\,s\leq t:\mathrm{TDB}(i,s)\text{ exists}\}$ denote the most recent time at or before $t$ when task $i$ was sampled (and its approximated regret stored). Then we define
$\mathrm{TaskDifficulty}(i,t)=\begin{cases}\mathrm{TDB}\bigl{(}i,\,s_{i}(t)\bigr{)},&\text{if }s_{i}(t)\text{ is finite},\\[4.0pt] 0,&\text{if }i\text{ has never been sampled before }t.\end{cases}$ (5)
This ensures that $\mathrm{TaskDifficulty}(i,t)$ always reflects the most recent approximated regret for task $i$, with larger values indicating greater remaining challenge.
Definition 3 (Co-Learnability). Beyond difficulty, we wish to capture how training on one task accelerates progress on others. Let $\mathcal{T}_{t+1}$ be the set of tasks replayed at time $t+1$. We define
$\mathrm{CoLearnability}(i,t)=\frac{1}{|\mathcal{T}_{t+1}|}\sum_{j\in\mathcal{T}_{t+1}}\bigl{[}\mathrm{TaskDifficulty}(j,t)-\mathrm{TaskDifficulty}(j,t+1)\bigr{]}$ (6)
i.e.,  the average reduction in difficulty of the replayed tasks induced by sampling $i$. Positive co-learnability indicates that visiting $i$ yields broad transfer benefits.
Proposition 1. Assuming $\mathcal{T}_{t+1}=\Lambda$ and defining
$Y_{j}\;=\;\mathrm{TaskDifficulty}(j,t+1)\;-\;\mathrm{TaskDifficulty}(j,t),$
the co-learnability of task $i$ at time $t$ coincides with the conditional expectation, that is
$\mathrm{CoLearnability}(i,t)=\mathbb{E}_{j\sim\mathcal{T}_{t+1}}\bigl{[}Y_{j}\mid T_{i}=1,\,t\bigr{]}.$
Proposition 1 confirms that our co-learnability definition precisely measures the mean one-step change in task difficulties, conditional on having replayed task i, when the next sampled task set covers the entire buffer.
Definition 4 (Task Priority). We combine difficulty and Co-Learnability into a scalar score, then pass it through a rank function, which replaces raw values by their relative ordering, to obtain robust sampling weights as in ACCEL [7]:
@@
where $\beta>0$ trades off challenge versus transfer. We sample tasks for replay proportionally to their computed priority scores, so that higher-priority tasks are more likely to be selected at each step. By ranking, we prevent outliers from dominating the curriculum and ensure a smooth sampling distribution.
Theorem 2 (Expected Staleness). Assume that at each replay step $s\geq 0$, tasks in buffer $\Lambda$ are sampled according to
$\Pr\bigl{[}i\text{ at }t_{0}+s\bigr{]}=\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)}\,,$
[…]
Then the expected number of steps $T=t_{i}-t_{0}$ until task $i$ is sampled satisfies
$\mathbb{E}[T]\;\leq\;\frac{|\Lambda|\,M_{0}}{\bigl{(}\mathrm{TaskPriority}(i,t_{0})\bigr{)}}\,.$
This guarantees that no task remains “stale” indefinitely: its expected waiting time scales inversely with its initial priority. While the boundedness and non-increasing requirements in Theorem 3.2 may initially seem restrictive, they follow naturally from the agent’s continual learning process. Indeed, Figure 10 shows an overall upward trend in returns during training, indicating that the agent’s performance steadily improves and, consequently, task priority scores tend to decrease over time.

### v3

@@ Abstract
Generalizing deep reinforcement learning agents to unseen environments remains a significant challenge. One promising solution is Unsupervised Environment Design (UED), a co‑evolutionary framework in which a teacher adaptively generates tasks with high learning potential, while a student learns a robust policy from this evolving curriculum. Existing UED methods typically measure learning potential via regret, the gap between optimal and current performance, approximated solely by value‑function loss. Building on these approaches, we introduce the transition-prediction error as an additional term in our regret approximation. To capture how training on one task affects performance on others, we further propose a lightweight metric called Co-Learnability. By combining these two measures, we present Transition‑aware Regret Approximation with Co‑learnability for Environment Design (TRACED). Empirical evaluations show that TRACED produces curricula that improve zero-shot generalization over strong baselines across multiple benchmarks. Ablation studies confirm that the transition-prediction error drives rapid complexity ramp‑up and that Co‑Learnability delivers additional gains when paired with the transition-prediction error. These results demonstrate how refined regret approximation and explicit modeling of task relationships can be leveraged for sample-efficient curriculum design in UED. https://geonwoo.me/traced/
@@ 1 Introduction
Deep reinforcement learning (RL) has achieved remarkable success in games, continuous control, and robotics (Sutton et al., 1998). Ideally, we want agents that generalize robustly to a broad range of unseen environments. However, hand‑crafting a training distribution that captures all real‑world variability is intractable, and agents often overfit even large training sets, performing poorly out‑of‑distribution (Kirk et al., 2023; Korkmaz, 2024).
@@
Unsupervised Environment Design (UED) tackles this by adapting the curriculum: a teacher module generates training tasks that challenge the student agent (Dennis et al., 2020). A popular class of UED methods measures task difficulty by regret, the difference between the optimal return and the agent’s achieved return, and uses this metric to guide curriculum design (Dennis et al., 2020; Jiang et al., 2021b; a; Parker-Holder et al., 2022; Azad et al., 2023b; Mediratta et al., 2023; Erlebach & Cook, 2024). Unfortunately, genuine regret requires knowing each environment’s optimal $Q^{*}$, which is infeasible in complex domains. Existing approaches, therefore, resort to coarse proxies such as Positive Value Loss (PVL) or maximum observed return (MaxMC) (Rutherford et al., 2024). In this paper, we refine regret estimation by augmenting PVL with a transition-prediction error term that captures how poorly a learned model predicts the environment’s dynamics. This combined signal provides a more faithful approximation of regret and a sharper basis for curriculum design.
@@
We further introduce a metric called Co‑Learnability to quantify how training on one task benefits others. For instance, consider three 100‑word vocabulary tasks, Spanish, English, and Japanese, whose transfer patterns differ: because Spanish and English share many cognates, learning Spanish accelerates lexical access and boosts English accuracy (Ramírez et al., 2013; Costa et al., 2000; Lemhöfer & Dijkstra, 2004), reflecting high Co‑Learnability; in contrast, Japanese is typologically distant (Chiswick & and, 2005), so gains from Japanese may transfer unefficiently to English, reflecting low Co‑Learnability. We present a lightweight estimator of Co‑Learnability that leverages observed changes in approximated regret, avoiding any additional modeling overhead in the UED loop.
@@
Figure 1: Task Priority Landscape. Task with high difficulty and high Co‑Learnability are scheduled with the highest priority in the curriculum.
@@
We propose TRACED (Transition-aware Regret Approximation with Co-Learnability for Environment Design), which combines refined regret and Co-Learnability to yield a task-priority landscape (Figure 1). We evaluate TRACED on procedurally generated MiniGrid (MG) (Chevalier-Boisvert et al., 2023) and BipedalWalker (BW) (Romac et al., 2021). In MG, we compare against Domain Randomization (DR) (Jakobi, 1997), PLR (Jiang et al., 2021b), ADD (Chung et al., 2024), and ACCEL (Parker-Holder et al., 2022) (the strongest baseline). In BW, we additionally include the state-of-the-art (SOTA) method CENIE (Teoh et al., 2024).
@@
TRACED surpasses all baselines in mean solved rate across 12 MiniGrid mazes, BipedalWalker, and even the extreme PerfectMaze variants. Ablation studies show that the transition-prediction error term accelerates curriculum ramp-up, while Co-Learnability provides additional gains when combined with our regret term. An analysis of curriculum evolution indicates that TRACED progressively increases task difficulty from easy to challenging. Taken together, these results chart a path toward more sample-efficient UED by coupling refined regret approximation with explicit modeling of task relationships.
@@ 2.1 Underspecified Partially Observable MDPs (UPOMDPs)
We model our environments as underspecified partially observable Markov decision processes (UPOMDPs) following Dennis et al. (2020). A UPOMDP is a tuple $\mathcal{M}=\langle A,O,\Theta,S,P_{0},P_{T},\mathcal{I},\mathcal{R},\gamma\rangle$, where $A$ is a finite set of actions, $S$ is a latent state space, and $O$ is an observation space. The observation function $\mathcal{I}:S\to\Delta(O)$ generates each observation $o_{t}$ given the true state $s_{t}$, the reward function $\mathcal{R}:S\times A\to\mathbb{R}$ and discount factor $\gamma\in[0,1)$ are shared across all levels. Crucially, $\Theta$ is a set of underspecified parameters that distinguish individual “levels”: for each $\theta\in\Theta$, the initial state is drawn from $P_{0}(\theta)\in\Delta(S)$ and transitions follow $P_{T}(s_{t+1}\mid s_{t},a_{t},\theta)$.
At each time step $t$, the agent observes $o_{t}\sim\mathcal{I}(s_{t})$ and selects an action $a_{t}$ according to a trajectory‑conditioned policy $\pi\bigl(a_{t}|o_{0},a_{0},\dots,o_{t}\bigr)$. For a fixed level $\theta$, the utility of policy $\pi$ is the expected discounted return $U_{\theta}(\pi)=\mathbb{E}\!\Bigl[\sum_{t=0}^{T}\gamma^{t}\,r_{t}\Bigr],r_{t}=\mathcal{R}(s_{t},a_{t})$ with the expectation taken over both the stochastic dynamics and the policy’s choices. We denote an optimal policy on level $\theta$ by $\pi^{\star}_{\theta}\in\arg\max_{\pi}U_{\theta}(\pi)$.
@@ 2.2 Unsupervised Environment Design (UED)
UED provides a series of levels with unknown parameters that are used to produce task environments automatically as a curriculum for the agent so as to efficiently train a single generalist policy $\pi_{\phi}$ across the entire parameter space $\Theta$ (Dennis et al., 2020). Recent UED methods maximize regret of the agent to generate a distribution of environments that guide effective learning (Jiang et al., 2021b; Parker-Holder et al., 2022; Teoh et al., 2024). The regret of policy $\pi$ is defined as the difference in expected reward between optimal policy $\pi^{*}$ and policy $\pi$, $Regret_{\theta}(\pi)=U_{\theta}(\pi^{*})-U_{\theta}(\pi)$.
Because the true optimal policy $\pi^{\star}$ is unknown in complex environments, existing methods approximate the instantaneous regret using proxy metrics. For example, PLR (Jiang et al., 2021a) evaluates two such proxies: Positive Value Loss (PVL) and Maximum Monte Carlo (MaxMC). PVL estimates the instantaneous regret as the average positive part of the Generalized Advantage Estimation (GAE)‐based Temporal Difference (TD) errors over an episode. PVL for an episode $\tau$ of length $T$ is defined as
@@
where $\gamma$ is the discount factor, $\lambda$ is the GAE coefficient and $\delta_{t}=r_{t}+\gamma\,V(s_{t+1})-V(s_{t})$ is the one‐step TD‐error at timestep $t$. MaxMC uses the highest undiscounted return observed on the task instead of a bootstrap target. Other criteria such as policy entropy, one‑step TD error, GAE, policy min‑margin, and policy least‑confidence have also been evaluated in curriculum learning contexts (Jiang et al., 2021b).
3 TRACED: Transition-aware Regret Approximation with Co-learnability for Environment Design
TRACED improves UED via (i) a regret approximation combining value and transition-prediction errors and (ii) a Co-Learnability measure. These yield a unified Task Priority score that governs new-task generation and replay sampling. We first present the motivating regret decomposition, then the curriculum derived from the approximated regret.
@@ 3.1 Regret Approximation via Transition Prediction Loss
Since regret quantifies the difficulty an agent experiences on a task, improving its approximation can yield more accurate difficulty estimates and, in turn, more effective curricula. We approximate the one-step regret at a state-action pair $(s,a)$ via the following decomposition:
$\mathrm{Regret}(s,a)$ $=V^{*}(s)-Q^{\pi}(s,a)$ (2)
$=V^{*}(s)-\hat{V}^{*}(s)+\hat{V}^{*}(s)-Q^{\pi}(s,a)$
$=\underbrace{V^{*}(s)-\hat{V}^{*}(s)}_{\text{(i) Value estimation error}}+\underbrace{r(s,a^{*})-r(s,a)}_{\text{(ii) Reward gap}}$
$\quad+\gamma\,\underbrace{\Bigl(\mathbb{E}_{s^{\prime\prime}\sim\hat{P}(\cdot\mid s,a^{*})}\!\bigl[\hat{V}^{*}(s^{\prime\prime})\bigr]-\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\!\bigl[V^{\pi}(s^{\prime})\bigr]\Bigr)}_{\text{(iii) Future value gap}}\,$,
Notation. A hat (e.g., $\hat{V}^{*}(s)$) indicates an empirical/learned estimator of the optimal value; $Q^{*},V^{*}$ are the optimal value functions; $Q^{\pi},V^{\pi}$ are the value functions under policy $\pi$; $P$ is the true transition kernel and $\hat{P}$ is a learned transition model; $\gamma$ is the discount factor.
Eq. 2 makes clear that Positive Value Loss (PVL), which evaluates only the accuracy of the empirical value estimator and thus corresponds to term (i), is insufficient as a proxy for regret. The future-value gap in term (iii) is influenced not only by value-function error but also by mismatch between the learned dynamics $\hat{P}$ and the true dynamics $P$. Motivated by this, we augment PVL with a transition-prediction error term, explicitly accounting for model-environment dynamics mismatch when approximating regret.
Definition 1 (Average Transition Prediction Loss). Train a transition dynamics estimator $f_{\phi}$, implemented as a recurrent model, to minimize a one‐step reconstruction loss $L_{\mathrm{trans}}(s_{t},a_{t})$ between the observed next state $s_{t+1}$ and the prediction $\hat{s}_{t+1}=f_{\phi}(s_{t},a_{t})$ . Detailed design choices for the transition model are provided in Appendix H.3. Over an episode $\tau=(s_{0},a_{0},\dots,s_{T})$, define
@@
where $\alpha>0$ balances value versus transition terms.
By explicitly incorporating both components, our regret approximation more faithfully captures task difficulty and, in turn, yields more effective curricula. For each sampled task instance $\tau$, we compute $\widehat{\mathrm{Regret}}(\tau)$ and append it to the task difficulty buffer (TDB), which is subsequently used to drive curriculum updates (Figure 2).
Figure 2: Task Difficulty Calculation Workflow. The Task Difficulty Buffer (TDB) records each task’s history of approximated regret. The agent interacts with sampled tasks to collect episode trajectories, which are stored in the Rollouts Buffer (RB). For each trajectory, we compute the Positive Value Loss (PVL) and the Average Transition‑Prediction Loss (ATPL). Their sum produces the updated task difficulty (approximated regret), which is appended to the TDB to refresh each sampled task’s stored difficulty.
@@ 3.2 Task Priority Construction
A central challenge in UED is to decide which tasks to present to the agent at each step (Hughes et al., 2024). Intuitively, we want to (1) focus on tasks that remain challenging, while (2) exploiting tasks whose training yields transfer to others. To this end, we introduce two complementary quantities, Task Difficulty and Co‑Learnability, and combine them into a single Task Priority score. In the following sections, time $t$ indexes the teacher’s curriculum-selection cycles (not environment time).
Definition 2 (Task Difficulty). Let $s_{i}(k)\;=\;\max\{\,s\leq k:\mathrm{TDB}(i,s)\text{ exists}\}$ denote the most recent time at or before $k$ when task $i$ was sampled, so that its approximated regret was stored at time $s_{i}(k)$. We define the task difficulty of task $i$ at time $k$ as
[…]
This ensures that $\mathrm{TaskDifficulty}(i,k)$ always reflects the most recent approximated regret for task $i$, with larger values indicating greater remaining challenge.
Definition 3 (Co‑Learnability). Beyond difficulty, we wish to capture how training on one task accelerates progress on others. Let $\mathcal{T}_{k+1}$ be the set of tasks replayed at time $k+1$. We define the Co-Learnability (CL) of task i at time k as
$\mathrm{CoLearnability}_{i}(k)=\frac{1}{|\mathcal{T}_{k+1}|}\sum_{j\in\mathcal{T}_{k+1}}\bigl[\mathrm{TaskDifficulty}(j,k)-\mathrm{TaskDifficulty}(j,k+1)\bigr]$ (6)
which measures the average reduction in the difficulty of replayed tasks when task $i$ is selected at time $k$. In principle, the true marginal contribution could be computed via Shapley values; however, due to computational constraints we approximate task $i$’s effect on reducing other tasks’ difficulty using this surrogate. A positive Co-Learnability value indicates that visiting $i$ yields transfer benefits.
Definition 4 (Task Priority). We combine difficulty and Co-Learnability into a scalar score and then apply a rank transform, which replaces raw values with their relative order (e.g., the largest value receives rank 1, the next largest rank 2, etc.), as in ACCEL (Parker-Holder et al., 2022):
@@
where $\beta>0$ trades off challenge versus transfer. At each step, we sample tasks inversely to their priority (e.g., $p(i\mid t)\propto 1/\mathrm{TaskPriority}(i,t)$), so lower ranks (higher priority) are selected more often. The rank transform prevents outliers from dominating the sampling distribution.